• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 00:02:06 on June14,2002

    Entry number 86009

    keyword=Shift Summary

    Lightning strikes started the shift. No beam for the first hour.

    17:15 beam back. Raster is off, optics target is in. Set beam position back to nominal (0,0) on bpmb. Setting up for bcm calibration. This one will do a half dozen measurements with 90 seconds of beam, followed by 90 seconds of no beam. Starting at 105 uA, drop current by 10 uA and do another measurements. Repeat all the down. On the way up, we go in coarser 20 uA steps.

    17:25 starting bcm calibration.

    17:35 MCC called. They need to steer beam in the injector (too many blm trips and the injector orbit had drifted) and will take beam away. Since the bcm calibration had just started, they will start at the beginning when they come back.

    21:30 Called MCC for update. Still steering up the injector.

    24:00 Still no beam.